Solve the Proportion for x: 21/x = 7/5
Answer by oscargut(2103)

Solve the Proportion for x: 21/x = 7/5

21 = 7x/5
21(5)=7x
x = (21/7)5
x = 3(5)
x = 15
Answer: x = 15
If you need more help with math questions/assignments/tests
You can ask me more using this e-mail (mthman@gmail.com) or you can visit my site:
http://mathshelp-online.com/
(Payment is using paypal)